XLib Hacking!

Student Information Processing Board (SIPB) Massachusetts Institute of Technology Room W20-557; x3-7788 IAP 1991

                          Chee Chew (web@athena.mit.edu)

Syllabus

1  The X Windows Model
2  X Windows Application Model
3  Real Programmers do it with lib
4  Mutilate a sample program
5  Your turn... Come up with something you want to write.

Resources

In addition to the SIPB (Room W20-557, phone x3-7788), the following resources are available:

Books

Athena Docs Athena does provide some documentation on Xlib programming, but it is NOT supported by OLC. You can get this documentation at Graphic Arts for a nominal fee.

The O’Reilly Books Are generally considered the best references for Xlib. Volume 1 is a manual which teaches how to program in Xlib, and Volume 2 is a very handy reference manual on all Xlib calls.

On-line Athena

Man Pages ARE YOUR FRIEND! type: man XOpenDisplay for example, and you will then get the manual page for XOpenDIsplay.

x11 locker has examples and sources for X, the server, clients, libraries, virtually everything you’d ever want to know about XLib, if you have the motivation to find it.